# Lessons on Datasets and Paradigms in Machine Learning for Symbolic Computation: A Case Study on CAD

### v1

@@ 1. Introduction
There has been a trend in recent years for programmers to train machine learning models to make these choices. Machine Learning (ML) refers to statistical techniques that learn rules from data. It has been well documented that ML has outperformed expert humans in a wide range of fields, driven forward by growth in computing power, new techniques, and methods, and an explosion in available training data. It has been suggested that ML can enhance the effectiveness of symbolic computation algorithms by replacing human-designed heuristics. This would allow non-expert users of symbolic computation to optimise algorithms to their application domain, and for mathematicians to focus on the theory with less distraction from such implementation details. Further, the emergence of Explainable AI techniques may allow for ML to provide suggestions and guidance on algorithm development, even for those who do not wish to incorporate ML into their final software [33].
@@ 1.2. Data augmentation
1.3. Our case study problem
@@ 1.4. Paper outline
In Section 5 we move on to reframe this problem in the the regression paradigm, comparing to classification and in Section 6 we present an alternative formulation that is more widely applicable. Finally, in Section 7 we share our conclusions obtained with these experiments.
@@ 2. Literature Review
ML has been improving multiple fields over the past two decades. In some safety-critical areas, such as automatic driving and medicine, ML models are not fully trusted, as they are not 100% accurate and an error could be fatal. However, they are still used to assist humans, such as when your car applies lane assist technology, or a doctor’s attention is drawn to particular cases. Similarly in mathematics, and thus symbolic computation, we usually require exact answers and so on the surface it may seem unlikely to use ML in computer algebra systems; except perhaps for drawing human attention to things of potential interest, as in [14], or where an ML-produced output can be easily checked for correctness, as in [29]. However, as discussed in the introduction, symbolic computation algorithms often contain choices that do not affect correctness but which have a great impact on the resources needed: such choices may be safely addressed by ML and are the focus of our study. We start by reviewing the prior literature for examples of this, paying particular attention to the learning paradigm utilised, before reviewing the literature for our case study problem.
@@ 2.1.1. Traditional classification paradigm
Most of the examples in the literature of ML to optimise symbolic computation take the form of supervised learning with the classification paradigm. Supervised ML means that we start with a dataset of labelled instances that we use to train the ML models. For example, a dataset of images where each image is labelled as either a cat or a dog is used to train a binary classifier to make that distinction. It is important that the data structure has some fixed features across the instances in the dataset, e.g. images of the same resolution or a representation as a fixed feature vector.
@@
There are further examples of supervised ML classification for computer algebra away from CAD in the literature. An early example was by Simpson et al. who used ML classifiers to pick from different algorithms to compute the resultant [36], substantially improving the runtimes in both Maple and Mathematica. ML classifiers to select the subalgorithm for a task have proved powerful in computational logic: where SATZilla was an early example to select a SAT-solver for a problem [39], and recently MachSMT did this for the far more diverse world of SMT [34].
@@ 2.1.3. Regression and reinforcement learning
As well as a move to regression instead of classification, this work also represents a shift from supervised to unsupervised learning. Peifer et al. did not measure the number of additional required for every possible $S$-pair in advance to create a dataset to learn from. Instead, their model made choices, got scored (the score being the number of additions), and amended its parameters in each iteration of the algorithm. This approach is usually known as reinforcement learning [^1] Another example of reinforcement learning is [27] which used it to learn the best pivot to use in Gaussian Elimination..
We note that [32] aimed to minimise the cost of each step, in the hope that if the cheapest decision is taken at each step, then the cost of the overall algorithm will be small. There is a risk with this approach that it leads to more $S$-pairs being generated overall, but [32] reported strong performance, with their reinforcement learning agent outperforming all human designed $S$-pair selection strategies. This motivates our own hypothesis that regression can allow for deeper learning than classification. However, unlike [32], we will not switch to reinforcement learning. We will directly compare supervised learning with classification and supervised learning with regression to see what, if any, performance difference there is.
@@ 2.2.3. Complexity
Davenport and Heinz proved that CAD generates a doubly exponential number of cells ($2^{2^{n}}$) in the worst case with respect to the number of variables ($n$) [13]. Brown and Davenport in [6] have shown that there is a family of sets of polynomials for which the right choice of variable ordering would yield a constant complexity and the wrong choice of variable ordering would result in a doubly exponential complexity. Numerous practical experiments have also shown the significant effect of the variable ordering choice in practice. This is the choice that we seek to optimise.
@@ 2.3.1. The Brown heuristic
It is not specified in [7] what to do if there is a third tie: our implementation randomly picks between the tied variables. Furthermore, whether the entire order of variables is chosen at once or one variable at a time is also not specified in [7].
2.3.2. Other human designed heuristics
@@ 3.1. Dataset
  • CAD timings: For each set of polynomials, CylindricalAlgebraicDecomposition from Maple’s Regular Chains Library [10] was timed to build the CAD for all six possible variable orderings. A time limit of 30 seconds was used and for the cases in which no variable ordering finished within the limit, the process was repeated with a time limit of 60 seconds.
  • Removal of over-complicated problems: Problems in which the construction of a CAD took more than 60 seconds for all six orderings were discarded, leaving 5599 problems.
  • Elimination of duplicates: Each problem has associated with it six CADs (one for each ordering) and thus a vector of six CAD cell counts. When we identified two problems for whom these vectors were identical then we discarded one problem. This happened extensively in the dataset where there are many problems that differ only by a single coefficient, having little effect on the geometry/topology of the problem. After this step, only 1019 problems were left. We note that this step is the main difference between our work and [22]: despite the latter using six times more data, they reported similar performance of the ML models trained. Aside from efficiency, this should avoid data leakage between testing and training.
@@ 3.2. Data embedding via features
There are various possible embeddings that ML models can take as input (e.g. images, text, numbers), but there is no ML model known to the authors that takes a set of polynomials directly as an input. In line with prior work on ML for the CAD variable ordering problem, we take a feature-engineering approach, representing each set of polynomials by a vector of floats defined from simple to compute metrics on the input such as those used by Brown and T1 above [25, 19, 11].
@@
    Extract the degree of the variable of each monomial. E.g. $I_{1,x_{1}}(f)=[3,2,2,1,0,0,0]$.
@@
    Represent with a 1 for each monomial containing the variable and with a 0 from each monomial not containing the variable. E.g. $I_{3,x_{1}}(f)=[1,1,1,1,0,0,0]$.
@@
The template that Florescu and England [19] proposed can be simplified to
$\operatorname*{operation}\limits_{p\in\mathcal{T}}\left(\operatorname*{operation}\limits_{M\in p}\left(\operatorname*{information}(M)\right)\right)$
where $\operatorname*{operation}$ is either a maximum, a sum or an average.
This allowed them to extract nine features per variable and information extraction. We note that [19] used $I_{1,x_{i}}(S),I_{3,x_{i}}(S)$ and $I_{4,x_{i}}(S)$ while we use $I_{1,x_{i}}(S),I_{2,x_{i}}(S)$ and $I_{3,x_{i}}(S)$ (which we found led to better performance). This obtains for us 27 features for each variable $x_{i}$. The number is reduced to 26 after $max(max(I_{3,x_{i}}(S)))$ is eliminated (which we do since it evaluated to $1$ for every instance in our dataset: every variable $x_{i}$ appears in every set of polynomials $S$ at least once). Then, motivated by Li et al. [30], for each variable we added another feature, the number of other variables that appear in the same polynomials as the variable under consideration, bringing the total back to 27 features per variable.
@@ 3.4. Train/test split
We will divide our dataset into five separate parts, known as “folds”. We will be careful to ensure that problems from the same source (the same directory of the SMT-LIB) end up in the same fold. This is to prevent “data leakage”: a situation in which a model is tested using the same or similar data to the data it was trained on. We will choose the hyperparameters and train models on four of these folds using the strategy being studied; then the models will be tested on the remaining fold. Repeating this until all folds have been used for testing, and taking the average to get final results. This process ensures that our testing is fair and that a model does not benefit from the test/train split, while still ensuring that the models are evaluated on data they have not seen during the selection of its hyperparameters or its training.
3.5. Hyperparameter tunning
In ML, “hyperparameters” are settings that govern the architecture and learning process of the model. Hyperparameters are set prior to training, in contract to model parameters which are optimised during training, e.g. the number of layers in a neural network or the number of branches in a decision tree. To find the best hyperparameters, we employ a technique called “Bayesian Optimisation” through the function BayesSearchCV in scikit-optimize. This function is more sophisticated than the more commonly used GridSearchCV which would evaluate all combinations of values for hyperparameters taken from discrete sets, and RandomisedSearchCV which randomly samples from such a grid. BayesSearchCV tries a fixed number of parameter samples from a specified distributions and intelligently navigates the hyperparameter space to select these.
@@ 3.6. ML models used
During our experiments, we will train a variety of models to ensure that the conclusions drawn generalise. The models we use are: K-Nearest Neighbours (KNN), Random Forest (RF), Support Vector Machine (SVM), Multi-Layer Perceptron (MLP), and Extreme Gradient Boosting (XGB). For the latter we use the implementation in the xgboost library and for the others the implementations in the sklearn library.
For training the models, we use the .fit() method in the sklearn library with default settings, maintaining a standardised process to fairly compare the changes under scrutiny.
@@ 3.7. Testing the models
Unless otherwise specified, the dataset that will be used to test the different strategies will be the balanced dataset described later in Section 4.3 to ensure the fairness of the comparisons.
In this paper, we aim to find a methodology to make good choices in symbolic computation. In particular, we focus on the case study of choosing good variable orderings. But what does “good” stand for in this context? To measure the complexity of a CAD we will use the time needed to generate it, but as it is impossible to obtain a model that guarantees choosing the optimal ordering every single time we must find ways to compare strategies that may be better in some instances but worse in others. We make use of the following metrics.
@@ 3.7.5. Metrics based on cell counts
An alternative set of metrics to those in Sections 3.7.2 -- 3.7.4 could be defined using a count of the number of cells of a CAD rather than the time taken to compute the CAD. In theory, such metrics would be implementation independent (although in practice, all CAD implementations differ on some of the underlying theory they implement). The cell count and the time taken should be closely correlated: we choose to focus on the time taken, as it is the measure of most value to users.
@@ 4. Improvements when using the Classification Paradigm
In this section, we will work with the supervised ML classification paradigm, most commonly used in the literature. First. in Section 4.1, the usual approach to creation of inputs for the ML model will be described. In Section 4.2 some problems in the created classification dataset will be identified, and in Section 4.3 some tools will be introduced to solve these problems. In Section 4.4 the misleading results that can be obtained if these issues are not detected are presented, and in the following subsections a series of improvements to the classification model are presented to overcome these difficulties. The first improvement solves the imbalanced problem in the training dataset, the second increases the size of the training dataset, and the last reduces the number of features describing an instance while preserving most of the information these features carry. These contributions were largely presented in the SC-Square 2023 workshop paper [15].
@@ 4.1.1. Creating the inputs
The task here is to represent a set of polynomials in three variables as a list of features. In Section 3.2 we described a methodology for creating a list of features, which we will denote by $f_{x}(\mathcal{T})$ for a given variable $x$ and a given set of polynomials $\mathcal{T}$.
To represent a three-variable set of polynomials $\mathcal{T}\in\mathbb{R}[x_{1},x_{2},x_{3}]$ to the classification model, we will therefore use the lists of features created with each variable appended to each other, placing the features related to the variable $x_{1}$ first, then those related to $x_{2}$ and finally those related to $x_{3}$. This list of features will be denoted by $F_{x_{1},x_{2},x_{3}}(\mathcal{T})=f_{x_{1}}(T)\,|\,f_{x_{2}}(\mathcal{T})\,|\,f_{x_{3}}(\mathcal{T})$.
@@ 4.1.2. Creating the labels
The labels correspond to the possible variable orderings that a CAD can use for a set of polynomials $\mathcal{T}\in\mathbb{R}[x_{1},x_{2},x_{3}]$. We will use as a label the variable ordering for which building a CAD took the shortest time. We will use the notation $\succ_{ijk}$ to refer to ordering $x_{i}\succ x_{j}\succ x_{k}$.

### v2

@@ 1. Introduction
There has been a trend in recent years for programmers to train machine learning models to make these choices. Machine Learning (ML) refers to statistical techniques that learn rules from data. It has been well documented that ML has outperformed expert humans in a wide range of fields, driven forward by growth in computing power, new techniques, and methods, and an explosion in available training data. It was first suggested a decade ago that ML might enhance the effectiveness of symbolic computation algorithms by replacing human-designed heuristics [26]. This could allow non-expert users of symbolic computation to optimise algorithms to their application domain, and for mathematicians to focus on the theory with less distraction from such implementation details. Further, the emergence of Explainable AI techniques may allow for ML to provide suggestions and guidance on algorithm development, even for those who do not wish to incorporate ML into their final software [34].
@@ 1.2. Data augmentation
1.3. Our case study problem: variable ordering for CAD
@@ 1.4. Paper outline
In Section 5 we move on to reframe this problem in the regression paradigm, comparing to classification, and in Section 6 we present an alternative formulation that is more widely applicable. Finally, in Section 7 we share our conclusions obtained with these experiments.
@@ 2. Literature Review
ML has been improving multiple fields over the past decades. In some safety-critical areas, such as automatic driving and medicine, ML models are not fully trusted, as they are not 100% accurate and an error could be fatal. However, they are still used to assist humans, such as when your car applies lane assist technology, or when a doctor’s attention is drawn to particular cases. Similarly in mathematics, and thus symbolic computation, we usually require exact answers and so on the surface it may seem unlikely to use ML in computer algebra systems; except perhaps for drawing human attention to things of potential interest, as in [15], or where an ML-produced output can be easily checked for correctness, as in [30]. However, as discussed in the introduction, symbolic computation algorithms often contain choices that do not affect correctness but which have a significant impact on the resources needed: such choices may be safely addressed by ML and are the focus of our study. We start by reviewing the prior literature for examples of this, paying particular attention to the learning paradigm utilised, before reviewing the literature for our case study problem.
@@ 2.1.1. Traditional classification paradigm
Most of the examples in the literature of ML to optimise symbolic computation take the form of supervised learning with the classification paradigm: we will survey those here. Supervised ML means that we start with a dataset of labelled instances that we use to train the ML models. For example, a dataset of images where each image is labelled as either a cat or a dog is used to train a binary classifier to make that distinction. It is important that the data structure has some fixed features across the instances in the dataset, e.g. images of the same resolution or a representation as a fixed feature vector.
@@
There are further examples of supervised ML classification for computer algebra away from CAD in the literature. An early example was by Simpson et al. who used ML classifiers to pick from different algorithms to compute the resultant [37], substantially improving the runtimes in both Maple and Mathematica. Recently, Barket et al. have experimented with ML to select the order in which to try 12 possible sub-algorithms for symbolic integration in Maple [2]. They experimented with both traditional LSTMs (Long-Short Term Memory neural networks) which process sequential data and a variant, Tree LSTMs, which processed the input as expression trees. They found the latter structure beneficial, and outperformed the standard Maple mechanism to select the sub-algorithm.
ML classifiers to select the sub-algorithm for a task have also proved powerful in computational logic: where SATZilla was an early example to select a SAT-solver for a problem [40], and recently MachSMT did this for the far more diverse world of SMT [35].
@@ 2.1.3. Regression and reinforcement learning
As well as a move to regression instead of classification, this work also represents a shift from supervised to unsupervised learning. Peifer et al. did not measure the number of additions required for every possible $S$-pair in advance to create a dataset to learn from. Instead, their model made choices, got scored (the score being the number of additions), and amended its parameters in each iteration of the algorithm. This approach is usually known as reinforcement learning [^1] Another example of reinforcement learning is [28] which used it to learn the best pivot to use in Gaussian Elimination..
We note that [33] aimed to minimise the cost of each step, in the hope that if the cheapest decision is taken at each step, then the cost of the overall algorithm will be small. There is a risk with this approach that it leads to more $S$-pairs being generated overall, but [33] reported strong performance, with their reinforcement learning agent outperforming all human-designed $S$-pair selection strategies. This motivates our own hypothesis that regression can allow for deeper learning than classification. However, unlike [33], we will not switch to reinforcement learning. We will directly compare supervised learning with classification and supervised learning with regression to see what, if any, performance difference there is.
Lastly, as the paper was being finalised, [27] was published with the first use of reinforcement learning to choose the variable ordering for CAD. Their methodology applied this approach with graph neural networks (using the same graph embedding discussed later in Section 2.3.2).
@@ 2.2.3. Complexity
Davenport and Heinz proved that CAD generates a doubly exponential number of cells ($2^{2^{n}}$) in the worst case with respect to the number of variables ($n$) [14]. Brown and Davenport in [7] have shown that there is a family of sets of polynomials for which the right choice of variable ordering would yield a constant complexity and the wrong choice of variable ordering would result in a doubly exponential complexity. Numerous practical experiments have also shown the significant effect of the variable ordering choice in practice, e.g. [18, 26]. This is the choice that we seek to optimise.
@@ 2.3.1. The Brown heuristic
It is not specified in [8] what to do if there is a third tie: our implementation randomly picks between the tied variables. Furthermore, whether the entire ordering of variables is chosen at once or one variable at a time taking into account intermediate projections is also not specified in [8].
2.3.2. Other human-designed heuristics
@@
Another recent heuristic was proposed by Li et al. (2021) [31]. This involved a graph structure associated with a polynomial set to capture sparsity information: the nodes of the graph are the variables in the polynomial set, with two nodes connected if they appear in the same polynomial. For example, the set of polynomials $S_{G}=\{x_{1}^{3}x_{2}-x_{1}+2,x_{2}^{4}-x_{3}\}$ would be associated with the graph in Figure 1. Their heuristic uses operations upon this graph to identify a variable ordering that seeks to preserve variable sparsity.
Figure 1. Graph associated to $S_{G}$.
@@ 3.1. Dataset
  • CAD timings: For each set of polynomials, CylindricalAlgebraicDecomposition from Maple’s Regular Chains Library [11] is timed to build the CAD for all six possible variable orderings. A time limit of 30 seconds is used and for the cases in which no variable ordering finished within the limit, the process was repeated with a time limit of 60 seconds.
  • Removal of over-complicated problems: Problems in which the construction of a CAD took more than 60 seconds for all six orderings are discarded, leaving 5599 problems.
  • Elimination of duplicates: Each problem has associated with it six CADs (one for each ordering) and thus a vector of six CAD cell counts. When we identify two problems for whom these vectors are identical then we discard one problem. This happens extensively in the dataset where there are many problems that differ only by a single coefficient, having little effect on the geometry/topology of the problem. After this step, only 1019 problems are left. We note that this step is the main difference between our work and [23]: despite the latter using six times more data, they reported similar performance of the ML models trained. Aside from efficiency, this should avoid data leakage between testing and training.
@@ 3.2. Data embedding via features
There are various possible embeddings that ML models can take as input (e.g. images, text, numbers), but there is no ML model known to the authors that takes a set of polynomials directly as an input. In line with prior work on ML for the CAD variable ordering problem, we take a feature-engineering approach, representing each set of polynomials by a vector of floats defined from simple-to-compute metrics on the input such as those used by Brown and T1 above [26, 20, 12].
@@
    Extract the degree of the variable of each monomial. E.g. $I_{1,x_{1}}(f)=[3,2,2,1,0,0,0]$. [^2] The order of monomials in a polynomial is not really relevant but is captured by the list data-structure: however, this order does not affect any of the operations performed upon the list later.
@@
    Represent with a '$1$' for each monomial containing the variable and with a '00' from each monomial not containing the variable. E.g. $I_{3,x_{1}}(f)=[1,1,1,1,0,0,0]$.
@@
The template that Florescu and England [20] proposed to extract features relative to the variable $x_{i}$ with respect to the set of polynomials $S$ can be simplified to
$\operatorname*{operation}\limits_{p\in S}\left(\operatorname*{operation}\limits_{M\in p}\left(I_{j,x_{i}}(S)\right)\right)$
where $\operatorname*{operation}$ is either a maximum, a sum, or an average, and where $j\in\{1,3,4\}$.
This allowed them to extract nine features per variable for each $j$. We note that [20] used $I_{1,x_{i}}(S),I_{3,x_{i}}(S)$ and $I_{4,x_{i}}(S)$ while we used $I_{1,x_{i}}(S),I_{2,x_{i}}(S)$ and $I_{3,x_{i}}(S)$ (which we found led to better performance). This obtains 27 features for each variable $x_{i}$. The number is reduced to 26 after $max(max(I_{3,x_{i}}(S)))$ is eliminated (which we do since it was evaluated to $1$ for every instance in our dataset: every variable $x_{i}$ appears in every set of polynomials $S$ at least once).
Finally, recall the work in [31] discussed in Section 2.3.2 earlier which utilised a graph embedding of the polynomial set. Inspired by their work, we add the degree of the variable in such a graph as a new feature, bringing the total back to 27 features per variable. In the example from Section 2.3.2 this new feature would evaluate to 1 for $x_{1}$ and $x_{3}$, and to 2 for $x_{2}$.
@@ 3.4. Train/test split
We divide our dataset into five separate parts, known as “folds”. We are careful to ensure that problems from the same source (the same directory of the SMT-LIB) end up in the same fold. This is to prevent “data leakage”: a situation in which a model is tested using the same or similar data to the data it was trained on. We choose the hyperparameters and train models on four of these folds using the strategy being studied; then the models will be tested on the remaining fold. Repeating this until all folds have been used for testing, and taking the average to get final results. This process ensures that our testing is fair and that a model does not benefit from the test/train split, while still ensuring that the models are evaluated on data they have not seen during the selection of its hyperparameters or its training.
3.5. Hyperparameter tuning
In ML, “hyperparameters” are settings that govern the architecture and learning process of the model. Hyperparameters are set prior to training, in contrast to model parameters which are optimised during training. Examples of hyperparameters would include the number of layers in a neural network or the number of branches in a decision tree.
To find the best hyperparameters, we employ a technique called “Bayesian Optimisation” through the function BayesSearchCV in scikit-optimize. This function is more sophisticated than the more commonly used GridSearchCV which would evaluate all combinations of values for hyperparameters taken from discrete sets, and RandomisedSearchCV which randomly samples from such a grid. BayesSearchCV tries a fixed number of parameter samples from specified distributions and intelligently navigates the hyperparameter space to select these.
@@ 3.6. ML models used
During our experiments, we train a variety of models to ensure that the conclusions drawn generalise. The models we use are: K-Nearest Neighbours (KNN), Random Forest (RF), Support Vector Machine (SVM), Multi-Layer Perceptron (MLP), and Extreme Gradient Boosting (XGB). For the latter we use the implementation in the xgboost library and for the others the implementations in the sklearn library. Models are trained using the .fit() method in the sklearn library with default settings, maintaining a standardised process to fairly compare the changes under scrutiny.
@@ 3.7. Testing the models
Unless otherwise specified, the dataset used to test the different strategies is the balanced dataset described later in Section 4.3 which best ensures the fairness of the comparisons.
In this paper, we aim to find a methodology to make good choices in symbolic computation. In particular, we focus on the case study of choosing good variable orderings. But what does “good” stand for in this context? To measure the complexity of a CAD we use the time needed to generate it, but as it is impossible to obtain a model that guarantees choosing the optimal ordering every single time we must find ways to compare strategies that may be better in some instances but worse in others. We make use of the following metrics.
@@ 3.7.4. Time markup
Over a dataset we define the markup as the average of the markups for all instances.
@@ 3.7.5. Metrics based on cell counts
An alternative set of metrics to those in Sections 3.7.2 -- 3.7.4 could be defined using a count of the number of cells of a CAD rather than the time taken to compute the CAD. In theory, such metrics would be implementation independent (although in practice, all CAD implementations differ on some of the underlying theory they implement). The cell count and the time taken should be lightly correlated (at least before any post-processing of cells): we choose to focus on the time related metrics, as it is the measure of most value to users.
@@ 4. Improvements when using the Classification Paradigm
In this section, we work with the supervised ML classification paradigm, most commonly used in the literature. First. in Section 4.1, the usual approach to the creation of inputs for the ML model is described. In Section 4.2 some problems in the created classification dataset is identified, and in Section 4.3 some tools are introduced to solve these problems. In Section 4.4 the misleading results that can be obtained if these issues are not detected are presented, and in the following subsections a series of improvements to the classification model are presented to overcome these difficulties. The first improvement solves the imbalanced problem in the training dataset, the second increases the size of the training dataset, and the last reduces the number of features describing an instance while preserving most of the information these features carry. These contributions were largely presented in the SC-Square 2023 workshop paper [16].
@@ 4.1.1. Creating the inputs
The task here is to represent a set of polynomials in three variables as a list of features. In Section 3.2 we described a methodology for creating a list of features, which we will denote by $f_{x}(S)$ for a given variable $x$ and a given set of polynomials $S$.
To represent a three-variable set of polynomials $S\in\mathbb{R}[x_{1},x_{2},x_{3}]$ to the classification model, we will therefore use the lists of features created with each variable appended to each other, placing the features related to the variable $x_{1}$ first, then those related to $x_{2}$ and finally those related to $x_{3}$.
@@ 4.1.2. Creating the labels
The labels correspond to the possible variable orderings that a CAD can use for a set of polynomials $S\in\mathbb{R}[x_{1},x_{2},x_{3}]$. We will use as a label the variable ordering for which building a CAD took the shortest time. We will use the notation $\succ_{ijk}$ to refer to ordering $x_{i}\succ x_{j}\succ x_{k}$.